# MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@ 4.2 Extensions to Local Gradient Methods
5 Convergence Analysis
Our analysis focuses on the i.i.d.  setting as shown in (1). We set $\eta_{m}\equiv\eta\in(0,1)$ for all $m\in\llbracket M\rrbracket$, typical in datacenter environments for distributed training [13], where workers access the entire data set. The analysis of heterogeneous data settings is deferred to future work. Our convergence analysis, following [69], does not assume bounded gradient variance across $\mathbb{R}^{d}$. Instead, our adaptive local batch size strategies ensure the expected strong growth condition [36, 5] at each local iterate $x_{k}^{m}$ for all workers $m\in\llbracket M\rrbracket$. This approach relaxes the standard assumptions of uniform boundedness and expected strong growth conditions [72], which are challenging to verify except in the interpolation regime [53].
Assumption 1 ($L$-Lipschitz smoothness).
Assumption 2 ($\mu$-strong convexity).
We derive the following convergence results for the proposed adaptive batch size strategies as follows. Their proofs are deferred to Appendix A.
Theorem 1 (Strongly convex; $\mu>0$).
$\mathbb{E}[F(x_{\mathsf{out}})]-F^{\star}=\widetilde{\mathscr{O}}\left(L(HM+\eta^{2})\exp\left\{-\frac{\mu K}{10L(HM+\eta^{2})}\right\}\left\lVert x_{0}-x^{\star}\right\rVert^{2}\right),$ (12)
Theorem 2 (Convex; $\mu=0$).
Theorem 3 (Nonconvex).
The theorems confirm that Local SGD, using the local norm test with exact variance, achieves linear convergence rates for both convex and nonconvex objectives. This scaling is proportional to the number of local gradient steps $H$ and workers $M$, akin to Local SGD under constant batch sizes and bounded gradient variance [69].
Remark 1 (Choice of $\eta$).
Convergence rates depend quadratically on $\eta$, which controls the likelihood and size of increased local batch sizes. Choosing smaller $\eta$ values accelerates convergence but causes quick batch size escalation, undermining adaptive batch sizing and widening the generalization gap.
6 Numerical Experiments
We evaluate the efficiency of the training (total steps or wall-clock time) and the generalization performance of our strategies through image classification tasks. We use a three-layer convolutional neural network on the MNIST dataset [46] and a ResNet-50 [28] on the CIFAR-10 dataset [42], with four workers ($M=4$ GPUs). The models are trained using Local Adam and Local SGD with momentum (SHB), synchronizing only the models every $H$ steps. Further details and results are in Appendix B.
6.1 Three-layer CNN on MNIST
We demonstrate our strategies using a three-layer CNN for MNIST classification. The training involves 12 million samples (200 epochs) with global and local batch sizes set to 60,000 and 15,000, respectively. This task uses a fixed learning rate of 0.0008 without any rate schedules. Results are in Table 1.
Table 1: Adaptive local batch size strategies for three-layer CNN on MNIST; no. of total steps (steps), wall-clock time (in minutes), average local batch sizes (bsz.), and top validation accuracy (acc.; in %)
Schedule $H=16$ $H=4$ $H=1$
steps time bsz. acc. steps time bsz. acc. steps time bsz. acc.
Constant 1456 5.01 2048 99.10 1464 9.44 2048 99.06 1464 25.58 2048 99.08
Constant 352 6.84 8192 98.80 364 8.19 8192 98.85 366 12.37 8192 98.94
Constant 192 10.59 15000 98.40 196 11.41 15000 98.43 199 13.88 15000 98.54
$\eta=0.25$ 213 11.35 13414 98.51 221 11.94 13427 98.55 223 14.50 13467 98.66
$\eta=0.375$ 224 11.85 13077 98.61 228 12.16 13031 98.58 228 14.86 13078 98.71
$\eta=0.50$ 229 11.85 12615 98.58 235 11.73 12648 98.60 235 14.42 12748 98.72
Figure 1: Validation accuracy and local batch sizes of Local Adam with adaptive batch size strategies ($H=16$).
Values of $\eta$.
Table 1 shows that adaptive local batch size strategies can match or exceed the validation accuracy of constant batch size strategies. Specifically, with 16 local gradient steps and $\eta=0.25$, the adaptive strategy achieves an average batch size similar to a constant batch size of 15,000 (see Figure 1), but with superior accuracy and 21 additional steps. This confirms that adaptive strategies can mitigate the generalization gap in large-batch training. Furthermore, Figure 2 indicates that these strategies reach a maximum local batch size of 15,000 later in the training, enhancing hardware utilization and efficiency. Similar trends are noted for $H=4$ and $H=1$.
Figure 2: Validation accuracy and local batch sizes of Local Adam with adaptive batch size strategies with $\eta\in\{0.25,0.375,0.5\}$ and $H\in\{16,4,1\}$.
Local gradient steps $H$.
Smaller $H$ increases training time due to more frequent communication, and contrary to expectations, does not always enhance generalization. As Table 1 demonstrates, with a local batch size of 2048, $H=1$ and $H=4$ have lower accuracy compared to $H=16$. This pattern is consistent with adaptive batch sizes, where $H=4$ underperforms $H=16$ at $\eta=0.375$. Moreover, while $H=1$ leads to longer training times, it achieves higher accuracy.
6.2 ResNet-50 on CIFAR-10
We scale our experiment to ResNet-50 for CIFAR-10, using a training budget of 12 million samples (200 epochs) and setting maximum global and local batch sizes of 60,000 and 15,000, respectively. Our learning rate schedules include a 10% linear warmup and cosine decay, with peak and minimum rates of 0.5 and 0.05, respectively.
Table 2: Adaptive local batch size strategies for ResNet-50 on CIFAR-10; no. of total steps (steps), wall-clock time (in minutes), average local batch sizes (bsz.), and top validation accuracy (acc.; in %)
Schedule $H=16$ $H=4$ $H=1$
steps time bsz. acc. steps time bsz. acc. steps time bsz. acc.
Constant 2432 27.16 1024 92.31 2440 34.00 1024 92.65 2441 60.51 1024 91.11
Constant 608 34.99 4096 63.50 608 37.10 4096 57.04 610 46.39 4096 40.40
Constant 304 34.31 8192 12.51 304 35.34 8192 16.61 305 40.01 8192 14.78
$\eta=0.8$ 576 90.73 4306 81.95 416 138.39 6001 75.09 241 50.67 10380 25.32
$\eta=0.85$ 512 92.41 4900 81.95 912 141.51 2742 86.89 433 144.75 5767 72.66
$\eta=0.9$ 2816 96.98 885 93.59 1982 126.38 1261 92.53 3940 111.92 634 93.75
Values of $\eta$.
Table 3 shows that adaptive local batch size strategies can achieve comparable or superior validation accuracy to constant batch sizes with similar averages. For instance, with $H=16$ and $\eta=0.80$, the adaptive strategy reaches an average batch size akin to a constant batch size of 4096 but with nearly 50% higher accuracy and 32 fewer steps, illustrating the effectiveness of adaptive strategies in bridging the generalization gap in large-batch training.
Figure 3: Validation accuracy and local batch sizes of Local SHB with adaptive batch size strategies ($H=16$).
Local gradient steps $H$.
We first examine constant batch sizes (Table 3 and Figure 4). Local gradient methods, as noted in [50, 58, 24, 25], may improve generalization compared to synchronized methods ($H=1$), although we employ SHB in synchronized models instead of standard SGD. For adaptive local batch sizes, smaller $H$ values increase batch sizes more rapidly (Table 3), supporting our theory that a smaller $\eta$ should be used with a larger $H$ to maintain convergence rates of $\mathscr{O}(HM+\eta^{2})$. Refer to Figures 3 and B.2 for further illustrations.
Figure 4: Validation accuracy and local batch sizes of Local SHB with adaptive batch size strategies with $\eta\in\{0.25,0.375,0.5\}$ and $H\in\{16,4,1\}$.
Number of workers $M$.
We conducted additional experiments with $M=8$. Compared to the 4-worker scenario, increasing the number of workers reduces training time but lowers average accuracy, indicating poorer generalization. Nonetheless, the observations regarding local gradient steps $H$ and values of $\eta$ remain consistent with the 4-worker case.
Figure 5: Validation accuracy and local batch sizes of Local SHB with adaptive batch size strategies ($H=16$) with 8 workers.
Table 3: Adaptive local batch size strategies for ResNet-50 on CIFAR-10 with 8 workers
Schedule $H=16$ $H=4$ $H=1$
steps time bsz. acc. steps time bsz. acc. steps time bsz. acc.
Constant 2432 12.06 512 92.06 2440 15.45 512 92.87 2441 29.09 512 92.41
Constant 1216 11.19 1024 88.88 1220 12.93 1024 87.87 1220 19.72 1024 86.80
Constant 608 10.72 2048 66.94 608 11.59 2048 71.32 610 15.03 2048 59.57
$\eta=0.8$ 213 11.55 5666 17.88 216 11.82 5655 15.98 218 12.29 5650 16.84
$\eta=0.85$ 220 12.22 5610 19.47 222 12.30 5570 22.06 222 12.66 5569 21.99
$\eta=0.9$ 261 15.25 5236 28.78 277 16.92 5052 33.07 278 18.26 4959 36.83
7 Concluding Remarks
This work introduces adaptive local batch size strategies for local gradient methods in data parallelism, focusing on reducing communication overheads by increasing local batch sizes independently. Although our analysis and experiments are limited to identical data settings, the algorithm is applicable to heterogeneous scenarios. Future work will explore convergence in diverse settings and federated learning [54, 33], including personalized approaches [27]. Extensions to asynchronous SGD [55, 41, 57, 45, 30] and hierarchical SGD [50, 8, 75] are also promising for improving communication efficiency and addressing worker heterogeneity.
Acknowledgments
The research of Han Liu is supported by NIH R01LM01372201, NSF RI 1840857, NSF TRIPOD 1740735, NSF DMS1454377-CAREER, NSF IIS 1546482, along with an Alfred P. Sloan Fellowship. The research of Mladen Kolar is supported in part by NSF ECCS-2216912. This research is supported in part through the computational resources and staff contributions from the Data Science Institute at the University of Chicago, through its AI + Science Research Funding Initiatives.
Appendix
Appendix A Proofs of Main Text
We present the omitted proofs of the main text, mainly by simplifying the technique in [69].
@@ Appendix A Proofs of Main Text
Let us define
$(\forall k\in\mathbb{N})\quad\overline{x}_{k+1}\coloneqq\overline{x}_{k}-\frac{\alpha_{k}}{M}\sum_{m=1}^{M}g_{k}^{m},$ (A.1)
where $g_{k}^{m}\coloneqq\nabla F_{\mathcal{B}_{k}^{m}}(x_{k}^{m})$ and $\overline{x}_{0}\coloneqq x_{0}$.
@@
Definition A.1 ($\tau$-slow sequence; Definition 10 in in [69]).
@@ A.1 Technical Lemmas
$\left\lVert\sum_{m=1}^{M}v_{m}\right\rVert\leqslant\sum_{m=1}^{M}\left\lVert v_{m}\right\rVert.$ (A.2)
Lemma A.2.
@@
$\left\lVert\sum_{m=1}^{M}v_{m}\right\rVert^{2}\leqslant M\sum_{m=1}^{M}\left\lVert w_{m}\right\rVert^{2}.$ (A.3)
Lemma A.3.
@@
$\left\lVert x+y\right\rVert^{2}\leqslant(1+\rho)\left\lVert a\right\rVert^{2}+(1+1/\rho)\left\lVert b\right\rVert^{2}.$ (A.4)
Lemma A.4 (Cauchy–Schwarz inequality).
@@
$2\langle x,y\rangle\leqslant\rho\left\lVert a\right\rVert^{2}+\rho^{-1}\left\lVert b\right\rVert^{2}.$ (A.5)
Lemma A.5.
@@
$(\forall x\in\mathbb{R}^{d})\quad\left\lVert\nabla F(x)\right\rVert^{2}\leqslant 2L(F(x)-F^{\star}).$ (A.6)
Lemma A.6 (Lemma 2 in [68]; Lemma 13 in [69]).
@@
Lemma A.7 (Lemma 14 in [69]).
@@ A.2 Proof of Theorems 1, 2 and 3
Lemma A.8 (Descent lemma for ($\mu$-strongly) convex objectives).
Suppose that Assumptions 1 and 2 hold with $\mu>0$. Let $(x_{k}^{m})_{k\in\mathbb{N},m\in\llbracket M\rrbracket}$ be the sequence of the iterates of Local SGD (2) with the (exact variance) local norm test (8) with $\eta_{m}\equiv\eta\in(0,1)$. If the sequence of learning rates $(\alpha_{k})_{k\in\mathbb{N}}$ satisfies $\alpha_{k}\leqslant M/(4L(M+\eta^{2}))$ for all $k\in\mathbb{N}$, then the sequence$(\overline{x}_{k})_{k\in\mathbb{N}}$ defined in (A.1) satisfies
@@
The proof is similar to that of Lemma 25 in [69]. We state it here for convenience and completeness. Expanding we have
@@
$\leqslant M\sum_{m=1}^{M}\left\lVert\nabla F(x_{k}^{m})\right\rVert^{2}+\eta^{2}\sum_{m=1}^{M}\left\lVert\nabla F(x_{k}^{m})\right\rVert^{2}$ by (A.3) and (8) (A.7)
[…]
@@
$-2\langle\nabla F(x_{k}^{m}),x_{k}^{m}-x^{\star}\rangle\leqslant-\mu\left\lVert x_{k}^{m}-x^{\star}\right\rVert^{2}-2\left(F(x_{k}^{m})-F^{\star}\right).$ (A.8)
By Cauchy–Schwarz inequality (A.5) with $\rho=2L$, we have
@@
[…]
By (A.4), we also have
$-\left\lVert x_{k}^{m}-x^{\star}\right\rVert^{2}\leqslant-\frac{1}{2}\left\lVert\overline{x}_{k}-x^{\star}\right\rVert^{2}+\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}.$ (A.10)
Thus, plugging in (A.6), (A.9) and (A.10), we obtain
@@
Lemma A.9 (Descent lemma for nonconvex objectives).
Suppose that Assumption 1 holds. Let $(x_{k}^{m})_{k\in\mathbb{N},m\in\llbracket M\rrbracket}$ be the sequence of the iterates of Local SGD (2) with the (exact variance) local norm test (8) with $\eta_{m}\equiv\eta\in(0,1)$. If the sequence of learning rates $(\alpha_{k})_{k\in\mathbb{N}}$ satisfies $\alpha_{k}\leqslant M/(2L(M+\eta^{2}))$ for all $k\in\mathbb{N}$, then the sequence$(\overline{x}_{k})_{k\in\mathbb{N}}$ defined in (A.1) satisfies
@@
The proof is similar to that of Lemma 26 in [69]. We state it here for convenience and completeness.
@@
$\leqslant F(\overline{x}_{k})-\frac{\alpha_{k}}{M}\sum_{m=1}^{M}\langle\nabla F(\overline{x}_{k}),\nabla F(x_{k}^{m})\rangle+\frac{L\alpha_{k}^{2}(M+\eta^{2})}{2M^{2}}\sum_{m=1}^{M}\left\lVert\nabla F(x_{k}^{m})\right\rVert^{2}$ by (A.7)
$=F(\overline{x}_{k})-\left(\frac{\alpha_{k}}{M}-\frac{L\alpha_{k}^{2}(M+\eta^{2})}{2M^{2}}\right)\sum_{m=1}^{M}\left\lVert\nabla F(x_{k}^{m})\right\rVert^{2}$
$\qquad+\frac{\alpha_{k}}{M}\sum_{m=1}^{M}\langle\nabla F(x_{k}^{m})-\nabla F(\overline{x}_{k}),\nabla F(x_{k}^{m})\rangle.$
By Cauchy–Schwarz inequality (A.5) with $\rho=1$ and Assumption 1, we have
@@
Lemma A.10.
Let $(x_{k}^{m})_{k\in\mathbb{N},m\in\llbracket M\rrbracket}$ be the sequence of the iterates of Local SGD (2) with the (exact variance) local norm test (8) with $\eta_{m}\equiv\eta\in(0,1)$, and $(\overline{x}_{k})_{k\in\mathbb{N}}$ be the sequence defined in (A.1). If the sequence of learning rates $(\alpha_{k})_{k\in\mathbb{N}}$ satisfies $\alpha_{k}\leqslant 1/(10L(HM+\eta^{2}))$ for all $k\in\mathbb{N}$ and $(\alpha_{k}^{2})_{k\in\mathbb{N}}$ is $H$-slow decreasing, then we have
$\frac{3L}{M}\sum_{m=1}^{M}\mathbb{E}\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}\leqslant\frac{1}{10LHM}\sum_{m=1}^{M}\sum_{h=0}^{H-1}\mathbb{E}\left\lVert\nabla F(x_{k-h}^{m})\right\rVert^{2}.$ (A.11)
@@
$\qquad\qquad+\frac{3}{M}\sum_{m=1}^{M}\sum_{h=0}^{H-1}\alpha_{k-h}^{2}\mathbb{E}\left\lVert g_{k-h}^{m}-\nabla F(x_{k-h}^{m})\right\rVert^{2}$ by (A.4) and (A.3)
@@
$\frac{1}{M}\sum_{m=1}^{M}\mathbb{E}\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}\leqslant\frac{3(H+\eta^{2})\alpha_{k}^{2}}{K}\sum_{m=1}^{M}\sum_{h=0}^{H-1}\mathbb{E}\left\lVert\nabla F(x_{k-h}^{m})\right\rVert^{2}.$ (A.12)
@@
$\leqslant\frac{1}{5LM}\sum_{m=1}^{M}\sum_{k=0}^{K}w_{k}\mathbb{E}\left\lVert\nabla F(x_{k-h}^{m})\right\rVert^{2}.$ (A.13)
@@
Using Lemma A.10 with $\mu>0$ and (A.6), we have
$\frac{3L}{M}\sum_{m=1}^{M}\sum_{k=0}^{K}w_{k}\mathbb{E}\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}\leqslant\frac{2}{5M}\sum_{m=1}^{M}\sum_{k=0}^{K}w_{k}\mathbb{E}[F(x_{k-h}^{m})-F^{\star}].$ (A.14)
Since $\alpha_{k}\equiv\alpha\leqslant 1/(10L(HM+\eta^{2}))\leqslant M/(4L(M+\eta^{2}))$, by Lemma A.8, we have
@@
Multiplying the weight $w_{k}$ and summing over $k\in\llbracket 0,K\rrbracket$, and by (A.14), we have
@@
$\frac{1}{MW_{K}}\sum_{m=1}^{M}\sum_{k=0}^{K}w_{k}\mathbb{E}[F(x_{k}^{m})-F^{\star}]\leqslant\frac{10}{W_{K}}\sum_{k=0}^{K}\left(\frac{w_{k}}{\alpha_{k}}\left(1-\frac{\mu\alpha_{k}}{2}\right)\mathbb{E}\left\lVert\overline{x}_{k}-x^{\star}\right\rVert^{2}-\frac{w_{k}}{\alpha_{k}}\mathbb{E}\left\lVert\overline{x}_{k+1}-x^{\star}\right\rVert^{2}\right).$ (A.15)
We observe that $1-\mu\alpha/2\geqslant 1-\mu/(20L(HM+\eta^{2}))\geqslant 1-1/(8HM)$, we choose $w_{k}=(1-\mu\alpha/2)^{-k-1}$ which are $2H$-slow increasing. Then by Lemma A.6, we obtain the desired result. ∎
@@
For the convex case with $\mu=0$, we simply use Lemma A.6 with $w_{k}\equiv 1$ for all $k\in\mathbb{N}$. ∎
@@
By Lemma A.9 and (A.6), and define $u_{k}\coloneqq 4\cdot\mathbb{E}[F(\overline{x}_{k})-F^{\star}]$, we have

### v2

@@ Appendix A Main Algorithm
Algorithm A.2 Adaptive Batch Size Schedules for Local Gradient Methods (Actual Implementation)
Sample the local batch $\mathcal{B}_{k,h}^{m}$ of size $b_{k}^{m}\coloneqq|\mathcal{B}_{k,h}^{m}|$ uniformly from $\mathcal{D}_{m}$
[…]
Compute the averaged batch gradient of all workers: $\overline{g}_{k,H}\coloneqq\frac{1}{M}\sum_{m=1}^{M}\nabla F_{\mathcal{B}_{k,H}^{m}}(x_{k,H}^{m})$triangleright all-reduce
Compute $\mathrm{Var}_{i\in\mathcal{B}_{k}}\left(\nabla f(x_{k}^{m};\xi_{i}^{m})\right)=b_{k}^{m}\cdot\frac{1}{M-1}\sum_{m=1}^{M}\left\lVert\nabla F_{\mathcal{B}_{k,H}^{m}}(x_{k,H}^{m})-\overline{g}_{k,H}\right\rVert^{2}$
Compute the norm test statistic $\mathsf{T}_{k}^{m}\coloneqq\left\lceil\left(\mathrm{Var}_{i\in\mathcal{B}_{k}}\left(\nabla f(x_{k,H}^{m};\xi_{i}^{m})\right)\right)\Big{/}\left(M\eta^{2}\left\lVert\overline{g}_{k,H}\right\rVert^{2}\right)\right\rceil$
Determine the next local batch sizes: $b_{k+1}^{m}=\max\{\mathsf{T}_{k}^{m},b_{k}^{m}\}$
$B\leftarrow B+HMb_{k}^{m}$
[…]
Appendix B Proofs of Main Text
We present the omitted proofs of the main text, mainly by simplifying the technique in Stich and Karimireddy (2020).
@@ Appendix B Proofs of Main Text
Let us define a sequence of (virtual) averaged iterates $(\overline{x}_{k})_{k\in\mathbb{N}}$, given by
$(\forall k\in\mathbb{N})\quad\overline{x}_{k+1}\coloneqq\overline{x}_{k}-\alpha_{k}\overline{g}_{k},$ (B.1)
where $\overline{g}_{k}\coloneqq\frac{1}{M}\sum_{m=1}^{M}g_{k}^{m}$ with $g_{k}^{m}\coloneqq\nabla F_{\mathcal{B}_{k}^{m}}(x_{k}^{m})$ and $\overline{x}_{0}\coloneqq x_{0}$.
@@
Definition B.1 ($\tau$-slow sequence; Definition 10 in Stich and Karimireddy, 2020).
@@ B.1 Technical Lemmas
$\left\lVert\sum_{m=1}^{M}v_{m}\right\rVert\leqslant\sum_{m=1}^{M}\left\lVert v_{m}\right\rVert.$ (B.2)
Lemma B.2.
@@
$\left\lVert\sum_{m=1}^{M}v_{m}\right\rVert^{2}\leqslant M\sum_{m=1}^{M}\left\lVert v_{m}\right\rVert^{2}.$ (B.3)
Lemma B.3.
@@
$\left\lVert x+y\right\rVert^{2}\leqslant(1+\rho)\left\lVert a\right\rVert^{2}+(1+1/\rho)\left\lVert b\right\rVert^{2}.$ (B.4)
Lemma B.4 (Cauchy–Schwarz inequality).
@@
$2\langle x,y\rangle\leqslant\rho\left\lVert a\right\rVert^{2}+\rho^{-1}\left\lVert b\right\rVert^{2}.$ (B.5)
Lemma B.5.
@@
$(\forall x\in\mathbb{R}^{d})\quad\left\lVert\nabla F(x)\right\rVert^{2}\leqslant 2L(F(x)-F^{\star}).$ (B.6)
Lemma B.6 (Lemma 2 in Stich, 2019b; Lemma 13 in Stich and Karimireddy, 2020).
@@
Lemma B.7 (Lemma 14 in Stich and Karimireddy, 2020).
@@ B.2 Proofs of Theorems 1, 2 and 3
Lemma B.8 (Descent lemma for ($\mu$-strongly) convex objectives).
Suppose that Assumptions 1 and 2 hold with $\mu>0$. Let $(x_{k}^{m})_{k\in\mathbb{N},m\in\llbracket M\rrbracket}$ be the sequence of the iterates of Local SGD (2) with the (exact variance) local norm test (9) with $\eta_{m}\equiv\eta\in(0,1)$. If the sequence of learning rates $(\alpha_{k})_{k\in\mathbb{N}}$ satisfies $\alpha_{k}\leqslant M/(4L(M+\eta^{2}))$ for all $k\in\mathbb{N}$, then the sequence $(\overline{x}_{k})_{k\in\mathbb{N}}$ defined in (B.1) satisfies
@@
The proof is similar to that of Lemma 25 in Stich and Karimireddy (2020). We state it here for convenience and completeness. Expanding we have
@@
$\leqslant M\sum_{m=1}^{M}\left\lVert\nabla F(x_{k}^{m})\right\rVert^{2}+\eta^{2}\sum_{m=1}^{M}\left\lVert\nabla F(x_{k}^{m})\right\rVert^{2}$ by (B.3) and (12) (B.7)
$\leqslant 2L(M+\eta^{2})\sum_{m=1}^{M}\left(F(x_{k}^{m})-F^{\star}\right)$ by (B.6) (B.8)
@@
$-2\langle\nabla F(x_{k}^{m}),x_{k}^{m}-x^{\star}\rangle\leqslant-\mu\left\lVert x_{k}^{m}-x^{\star}\right\rVert^{2}-2\left(F(x_{k}^{m})-F^{\star}\right).$ (B.9)
By Cauchy–Schwarz inequality (B.5) with $\rho=2L$, we have
@@
$\leqslant 2L\left\lVert x_{k}^{m}-x^{\star}\right\rVert^{2}+F(x_{k}^{m})-F^{\star}$ by (B.6) (B.10)
By (B.4), we also have
$-\left\lVert x_{k}^{m}-x^{\star}\right\rVert^{2}\leqslant-\frac{1}{2}\left\lVert\overline{x}_{k}-x^{\star}\right\rVert^{2}+\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}.$ (B.11)
Thus, plugging in (B.8), (B.10) and (B.11), we obtain
@@
Lemma B.9 (Descent lemma for nonconvex objectives).
Suppose that Assumption 1 holds. Let $(x_{k}^{m})_{k\in\mathbb{N},m\in\llbracket M\rrbracket}$ be the sequence of the iterates of Local SGD (2) with the (exact variance) local norm test (9) with $\eta_{m}\equiv\eta\in(0,1)$. If the sequence of learning rates $(\alpha_{k})_{k\in\mathbb{N}}$ satisfies $\alpha_{k}\leqslant M/(2L(M+\eta^{2}))$ for all $k\in\mathbb{N}$, then the sequence $(\overline{x}_{k})_{k\in\mathbb{N}}$ defined in (B.1) satisfies
@@
The proof is similar to that of Lemma 26 in Stich and Karimireddy (2020). We state it here for convenience and completeness.
@@
$\leqslant F(\overline{x}_{k})-\frac{\alpha_{k}}{M}\sum_{m=1}^{M}\langle\nabla F(\overline{x}_{k}),\nabla F(x_{k}^{m})\rangle+\frac{L\alpha_{k}^{2}(M+\eta^{2})}{2M^{2}}\sum_{m=1}^{M}\left\lVert\nabla F(x_{k}^{m})\right\rVert^{2}$ by (B.7)
$=F(\overline{x}_{k})-\left(\frac{\alpha_{k}}{M}-\frac{L\alpha_{k}^{2}(M+\eta^{2})}{2M^{2}}\right)\sum_{m=1}^{M}\left\lVert\nabla F(x_{k}^{m})\right\rVert^{2}+\frac{\alpha_{k}}{M}\sum_{m=1}^{M}\langle\nabla F(x_{k}^{m})-\nabla F(\overline{x}_{k}),\nabla F(x_{k}^{m})\rangle.$
By Cauchy–Schwarz inequality (B.5) with $\rho=1$ and Assumption 1, we have
@@
Lemma B.10.
Let $(x_{k}^{m})_{k\in\mathbb{N},m\in\llbracket M\rrbracket}$ be the sequence of the iterates of Local SGD (2) with the (exact variance) local norm test (9) with $\eta_{m}\equiv\eta\in(0,1)$, and $(\overline{x}_{k})_{k\in\mathbb{N}}$ be the sequence defined in (B.1). If the sequence of learning rates $(\alpha_{k})_{k\in\mathbb{N}}$ satisfies $\alpha_{k}\leqslant 1/(10L(HM+\eta^{2}))$ for all $k\in\mathbb{N}$ and $(\alpha_{k}^{2})_{k\in\mathbb{N}}$ is $H$-slow decreasing, then we have
$\frac{3L}{M}\sum_{m=1}^{M}\mathbb{E}\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}\leqslant\frac{1}{10LHM}\sum_{m=1}^{M}\sum_{h=0}^{H-1}\mathbb{E}\left\lVert\nabla F(x_{k-h}^{m})\right\rVert^{2}.$ (B.12)
@@
$\qquad\qquad+\frac{3}{M}\sum_{m=1}^{M}\sum_{h=0}^{H-1}\alpha_{k-h}^{2}\mathbb{E}\left\lVert g_{k-h}^{m}-\nabla F(x_{k-h}^{m})\right\rVert^{2}$ by (B.4) and (B.3)
@@
$\frac{1}{M}\sum_{m=1}^{M}\mathbb{E}\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}\leqslant\frac{3(H+\eta^{2})\alpha_{k}^{2}}{K}\sum_{m=1}^{M}\sum_{h=0}^{H-1}\mathbb{E}\left\lVert\nabla F(x_{k-h}^{m})\right\rVert^{2}.$ (B.13)
@@
$\leqslant\frac{1}{5LM}\sum_{m=1}^{M}\sum_{k=0}^{K}w_{k}\mathbb{E}\left\lVert\nabla F(x_{k-h}^{m})\right\rVert^{2}.$ (B.14)
@@
Using Lemma B.10 with $\mu>0$ and (B.6), we have
$\frac{3L}{M}\sum_{m=1}^{M}\sum_{k=0}^{K}w_{k}\mathbb{E}\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}\leqslant\frac{2}{5M}\sum_{m=1}^{M}\sum_{k=0}^{K}w_{k}\mathbb{E}[F(x_{k-h}^{m})-F^{\star}].$ (B.15)
Since $\alpha_{k}\equiv\alpha\leqslant 1/(10L(HM+\eta^{2}))\leqslant M/(4L(M+\eta^{2}))$, by Lemma B.8, we have
@@
Multiplying the weight $w_{k}$ and summing over $k\in\llbracket 0,K\rrbracket$, and by (B.15), we have
@@
$\frac{1}{MW_{K}}\sum_{m=1}^{M}\sum_{k=0}^{K}w_{k}\mathbb{E}[F(x_{k}^{m})-F^{\star}]\leqslant\frac{10}{W_{K}}\sum_{k=0}^{K}\left(\frac{w_{k}}{\alpha_{k}}\left(1-\frac{\mu\alpha_{k}}{2}\right)\mathbb{E}\left\lVert\overline{x}_{k}-x^{\star}\right\rVert^{2}-\frac{w_{k}}{\alpha_{k}}\mathbb{E}\left\lVert\overline{x}_{k+1}-x^{\star}\right\rVert^{2}\right).$ (B.16)
We observe that $1-\mu\alpha/2\geqslant 1-\mu/(20L(HM+\eta^{2}))\geqslant 1-1/(8HM)$, we choose $w_{k}=(1-\mu\alpha/2)^{-k-1}$ which are $2H$-slow increasing. Then by Lemma B.6, we obtain the desired result. ∎
@@
For the convex case with $\mu=0$, we simply use Lemma B.6 with $w_{k}\equiv 1$ for all $k\in\mathbb{N}$. ∎
@@
By Lemma B.9 and (B.6), and define $u_{k}\coloneqq 4\cdot\mathbb{E}[F(\overline{x}_{k})-F^{\star}]$, we have